Prove that for any 0 < x < 1, the series (1−x)+(x²−x³)+... converges and show that it converges to 1/(1+x)


1
Expert's answer
2021-02-24T06:36:57-0500

Prove that for any 0 < x < 1, the series (1−x)+(x²−x³)+... converges and show that it converges to 1/(1+x)

Solution:

"\displaystyle(1−x)+(x^2−x^3)+(x^4−x^5)+...=\sum_{\mathclap{n=0}} ^{\mathclap{\infty}}(1-x)x^{2n}"

"\displaystyle\sum_{\mathclap{n=0}} ^{\mathclap{\infty}}(1-x)x^{2n}=\displaystyle\sum_{\mathclap{n=0}} ^{\mathclap{\infty}}(1-x)({x^2})^n"

"\displaystyle\sum_{\mathclap{n=0}} ^{\mathclap{\infty}}(1-x)({x^2})^n" is an infinite geometric series. It is an infinite series whose successive terms have a common ratio "x^2" . Such a series converges if and only if the absolute value of the common ratio is less than one ("|x^2| < 1" ).

Since:

"0 < x < 1"

then:

"0 < x^2 < 1" and the convergence condition is satisfied.

The sum of the series can be computed from the finite sum formula:

"\displaystyle\sum_{\mathclap{n=0}} ^{\mathclap{\infty}}(1-x)({x^2})^n=\lim_{\mathclap{k\to\infty}}\displaystyle\sum_{\mathclap{n=0}} ^{\mathclap{k}}(1-x)({x^2})^n="

"\displaystyle\lim_{\mathclap{k\to\infty}}\frac{(1-x)(1-(x^2)^{k+1})}{1-x^{2}}=\frac{1-x}{1-x^2}="

"\displaystyle\frac{1-x}{(1-x)(1+x)}=\frac{1}{1+x}"